Need a detailed explanation to question 2. I have the answer.

1. Suppose there are two firms, Boors and Cudweiser, each selling identical-tasting nonalcoholic beer. Consumers of this beer have no brand loyalty so market demand can be expressed as P = 5 − .001(QB + QC). Boors’ marginal revenue function can be written MR = 5 − .001(2QB + QC) and symmetrically for Cudweiser. Boors operates with out-of-date technology and has constant cost of $2 per unit , whereas Cudweiser has constant cost of $1 per unit. Assuming the firms behave as Cournot competitors, Boor’s best-response function is

a. QB = 2,000 − .5QC b. QB = 1,500 − .5QC c. QC = 2,000 − .5QB d. QC = 1,500 − .5QB

2. Consider the same market for nonalcoholic beer as in the previous question. Cudweiser’s response function is

a. QB = 2,000 − .5QC b. QB = 1,500 − .5QC c. QC = 2,000 − .5QB d. QC = 1,500 − .5QB
Answer #1

(2) (c)

For Cudweiser,

TR = P x QC = 5QC - 0.001QBQC - 0.001QC2

MR = TR/QC = 5 - 0.001QB - 0.002QC

setting MR = MC,

5 - 0.001QB - 0.002QC = 1

0.001QB + 0.002QC = 4

QB + 2QC = 4,000

2QC = 4,000 - QB

QC = 2,000 - 0.5QB (Best response function)